X is a random variable such that P(x=n)=Pn n=1,2,...



GP(x)=pgf of x.



Define P(x≤n)=q,and obtain the pgf of qn in terms of GP(x)

1
Expert's answer
2021-12-14T13:11:35-0500

The probability generating function (PGF) of X is GX(s) = E(sX), for all s ∈ R for which the sum converges:


"G_{xn}(s)=\displaystyle\sum_{k=1}^{n} s^kP(x=n)=\displaystyle\sum_{k=1}^{n} s^kP_n"


"G_{qn}(s)= sP(x=1)+\displaystyle\sum_{k=1}^{2} s^2P(x=2)+...+\displaystyle\sum_{k=1}^{n} s^nP(x=n)"


"G_{qn}(s)=\displaystyle\sum_{k=1}^{n}G_{xn}(s)"